The mmolality of 1M solution of sodium nitrate is 0.8mol/kg. Determine the density of the solution.

Grade:11

1 Answers

noogler
8 years ago
as the solution is same no. of moles ofsoluteis samein boththe cases
M=n/V & m=n/W
Density=W/V=M/m=1/0.8=1.25 kg/lit
 
 
aprrove if ans. is right